OpenStudy (anonymous):

What is 6+7x45(6+23)?

OpenStudy (karatechopper):

is that a 7x or a 7 times 45?

OpenStudy (anonymous):

7 times

OpenStudy (karatechopper):

Oohhh... ok! Try PEMDAS on that!

OpenStudy (anonymous):

I already did and couldn't do it. Please help me :(

OpenStudy (anonymous):

6+7x45(6+23) = 6+7.45(29) = 6+7.(45x29) = 6 + (7x45x29) =?

OpenStudy (ja1):

PEMDAS

OpenStudy (anonymous):

Thanks Daus! :D

OpenStudy (ja1):

Parenthesis Exponents Multiplication Division Addition Subtraction

OpenStudy (ja1):

Do the parenthesis first

OpenStudy (ja1):

\[(6+23) = ?\]

OpenStudy (ja1):

And @dauspex that's really not the most common or easy way to do this problem tbh

OpenStudy (anonymous):

@JA1 why do you say so? the first step was to remove the brackets, then i basically showed him that for a sector of 3 multiplied terms, it does not matter which order of multiplication is done

OpenStudy (ja1):

Yes but that is the way anyone with experience doing this type of math would do it, but for someone just learning it is best to stick to PEMDAS method like so: \[6+7\times45(6+23)\] \[6+7\times45\times29\] Now you multiply from left to right: \[6 + 315 \times 29\] \[6 + 9135\] Now "Addition": \[6141\] Doing it this way is following the PEMDAS method more directly so it is easier for him to learn :) @dauspex

OpenStudy (anonymous):

I get what you mean, good feedback though, we tend often to forget haha

OpenStudy (ja1):

It's fine xD you did good though, I just like to keep the perspective of a user that doesn't know how to do this :)

OpenStudy (anonymous):

Thanks for solving it in your argument.

OpenStudy (ja1):

No prob :) and welcome to OpenStudy ^,^